8 75° 40° 1. Given the information above: (1 pt) Which cannon ball will move through the air further? a. A b. B c. They will have the same horizontal displacement 、V.-loom, sama initial speed
Answer #1

The maximum horizontal distnace travelled by a projectile is given by v2sin2\theta /g .Therefore, horizontal distance is proportinal to sin2\theta

Given \theta for A is 75,sin2\theta = sin 150=0.5

\theta for B is 40, sin 2\theta=sin80=0.98

Therefore, Cannon B will move further because 0.98 is greater than 0.5

Correct option is b